# Backbone probability of planar Brownian motion

### v1

@@ Abstract
Motivated by critical planar percolation, we investigate a “backbone” event of planar Brownian motion, i.e. the existence of two disjoint subpaths on the Brownian trajectory connecting the $\varepsilon$-neighborhood of the starting point to a macroscopic distance. We show that the probability of this event is $(\log|\log\varepsilon|)^{-1}$ up to a multiplicative constant.
@@ 1.1 Overview and the main result
Note that in the critical percolation, one can also define the monochromatic arm exponents corresponding to that there exist $n$ disjoint arms of the same color joining two boundaries of an annulus. Based on the relation between Brownian motion and critical percolation, it is natural to explore the analog of such monochromatic arm exponents in the context of Brownian motion. Recently, [NQSZ24] derives the exact value of percolation monochromatic 2-arm exponent, namely the backbone exponent. In this paper, we investigate its Brownian counterpart, and find that such Brownian backbone probability indeed has an iterated logarithmic decay (so the “Brownian backbone exponent” is equal to 0). This also shows a big difference between Brownian motion and critical percolation, although their outer boundaries (or hulls) are the same.
@@
There exist constants $C_{1},C_{2}>0$ such that for any $\varepsilon\in(0,\frac{1}{10})$,
$\frac{C_{1}}{\log|\log\varepsilon|}\leq\mathbbm{P}[\mathsf{Bac}_{\varepsilon}]\leq\frac{C_{2}}{\log|\log\varepsilon|}.$ (1.2)
@@
The same result also holds when $\frac{1}{2}\mathbbm{S}^{1}$ in (1.1) is changed to any fixed $r\mathbbm{S}^{1}$ with $r\in(0,1)$ (and for sufficiently small $\varepsilon$), except that the corresponding constants in (1.2) will depend on $r$.
Note that the backbone event $\mathsf{Bac}_{\varepsilon}$ is closely related to the spatial distribution of the cut points of $B[0,\tau_{\mathbbm{D}}]$. Indeed, our proof of Theorem 1.1 is based on our recent paper with Fu and Sun [CFSX25], from which we can define a layer structure for the cut points of $B[0,\tau_{\mathbbm{D}}]$ and then solve it explicitly; see Section 3. We mention that [CFSX25] heavily relies on the connection between planar Brownian motion and $\operatorname{SLE}_{8/3}$, and especially their coupling with Liouville quantum gravity (LQG). It would be quite interesting to see if there is a derivation of Theorem 1.1 without relying on LQG.
@@ 1.2 Outlook and discussions
  • In our forthcoming work [CX25], we will extend the result of this paper to the Brownian loop soup cases. Let $(B_{t})_{0\leq t\leq\tau_{\mathbbm{D}}}$ be as before, and let $\mathcal{L}$ be an independent Brownian loop soup on $\mathbbm{D}$ with intensity $\frac{c}{2}$ for $c\in(0,1]$. For a subset $A\subset\overline{\mathbbm{D}}$, denote $\mathcal{C}(A)$ to be the union of $A$ and all loop-soup clusters in $\mathcal{L}$ intersecting with $A$, and let $H:=\mathcal{C}(B[0,\tau_{\mathbbm{D}}])$. However, unlike the Brownian motion case, there is a positive probability that 0 itself can have two disjoint subpaths (except on 0) on $H$ connected to $\frac{1}{2}\mathbbm{S}^{1}$. Moreover, in [CX25] we will show the following
@@
Organization of the paper. In Section 2, we review the explicit relation between the laws of conformal radii and moduli for $\operatorname{SLE}_{8/3}$-type loops. Such relation was implicitly established in [ARS25b], and we will provide its proof in Appendix A for completeness. Then in Section 3, we define a layer structure for the cut points of $B[0,\tau_{\mathbbm{D}}]$ and solve it explicitly, based on the results of [CFSX25] and Section 2. Finally, we finish the proof of Theorem 1.1 in Section 4.
@@
Acknowledgment. We are grateful to Xin Sun for many helpful discussions and suggestions on the early draft of this paper. G.C. and Z.X. were partially supported by National Key R&D Program of China (No. 2023YFA1010700). G.C. was partially supported by National Key R&D Program of China (No. 2021YFA1002700).
@@ 2 Conformal radii of $\operatorname{SLE}_{8/3}$-type loops
3 The exact laws of the conformal radii
@@
Let $(t_{i})_{i\geq 1}$ be the collection of cut time $t$’s of $B[0,\tau_{\mathbbm{D}}]$ such that the outer boundary of $B[0,t]$ is a simple loop, ranked in decreasing order. Note that a.s. $\lim_{i\to\infty}t_{i}=0$. For each $i\geq 1$, define $s_{i}$ to be the last cut time before $t_{i}$ (thus $t_{i}>s_{i}>t_{i+1}$). Let $\ell_{i}$ and $\widetilde{\ell}_{i}$ be the outer boundary of $B[0,t_{i}]$ and the boundary of the connected component of $\mathbbm{D}\setminus B[s_{i},t_{i}]$ containing the origin, respectively.
@@
    Let $g:D(\ell_{1})\to\mathbbm{D}$ be the conformal map such that $g(0)=0$ and $g(B_{t_{1}})=1$. Then $g(\widetilde{\ell}_{1})$ is independent of $\ell_{1}$. In particular, $\mathrm{CR}(\rho,0)$ is independent of $\ell_{1}$.
@@
    Let $\widetilde{g}:D(\widetilde{\ell}_{1})\to\mathbbm{D}$ be the conformal map such that $\widetilde{g}(0)=0$ and $\widetilde{g}(B_{s_{1}})=1$. Then $(\widetilde{g}(B_{t}))_{t\in[0,s_{1}]}$ is independent of $(B_{t})_{t\in[s_{1},\tau_{\mathbbm{D}}]}$, and has the same law (up to a time change) as $(B_{t})_{t\in[0,\tau_{\mathbbm{D}}]}$ conditioned on $B_{\tau_{\mathbbm{D}}}=1$.
As a corollary, if we set $\widetilde{\ell}_{0}=\mathbbm{S}^{1}$, then $\left(\frac{\mathrm{CR}(\widetilde{\ell}_{n},0)}{\mathrm{CR}(\ell_{n},0)}\right)_{n\geq 1}$ and $\left(\frac{\mathrm{CR}(\ell_{n+1},0)}{\mathrm{CR}(\widetilde{\ell}_{n},0)}\right)_{n\geq 0}$ are two i.i.d. r.v. sequences, and mutually independent of each other.
@@
We first prove (ii). Due to rotational invariance, it suffices to consider the case when conditioning on $B_{\tau_{\mathbbm{D}}}=1$. Let $(Y_{t})_{t\geq 0}$ be the time reversal of $(B_{t})_{t\in[0,\tau_{\mathbbm{D}}]}$ conditioned on $B_{\tau_{\mathbbm{D}}}=1$. Note that $Y_{t}$ is indeed a Brownian excursion in $\mathbbm{D}$ from 1 to 0. Let $(\mathcal{F}_{t})_{t\geq 0}$ be the natural filtration of $(Y_{t})_{t\geq 0}$, and $(\mathcal{G}_{t})_{t\geq 0}$ be the cut time filtration defined in [Vir03, Section 3] (i.e. $\mathcal{G}_{t}$ is the $\sigma$-field generated by $\mathcal{F}_{t}$ and the set of all cut times of $(Y_{t})_{t\geq 0}$ before time $t$). By [Vir03, Proposition 14], if $\tau$ is a $(\mathcal{G}_{t})_{t\geq 0}$-stopping time such that $\tau$ is also a cut time of $(Y_{t})_{t\geq 0}$, then the conditional law of $(Y_{\tau+u})_{u\geq 0}$ given $\mathcal{G}_{\tau}$ is a Brownian excursion from $Y_{\tau}$ to 0 in the connected component of $\mathbbm{D}\backslash Y[0,\tau]$ containing 0. Now we consider $s_{1}^{\prime}:=\tau_{\mathbbm{D}}-s_{1}$. Note that $s_{1}^{\prime}$ is the first cut time of $(Y_{t})_{t\geq 0}$ such that $Y[0,s^{\prime}_{1}]$ disconnects the origin from $\mathbbm{S}^{1}$, thus is a $(\mathcal{G}_{t})_{t\geq 0}$-stopping time. Therefore, the conditional law of $(\widetilde{g}(Y_{s_{1}^{\prime}+u}))_{u\geq 0}$ given $\mathcal{G}_{s_{1}^{\prime}}$ is an independent Brownian excursion (up to a time change) in $\mathbbm{D}$ from $1$ to 0. By taking the time reversal of $(Y_{t})_{t\geq 0}$, we obtain the result.
Then we prove (i). Let $A\subset\overline{\mathbbm{D}}$ be relatively compact such that $\A\mathbbm{D}\backslash$ A is simply connected and contains 0. Note that the law of $(B_{t})_{t\in[0,\tau_{\mathbbm{D}}]}$ conditioned on the event $B[0,\tau_{\mathbbm{D}}]\cap A=\emptyset$ is the same as the Brownian motion (up to a time change) in $\A\mathbbm{D}\backslash$ A starting from 0 and conditioned on exiting $\A\mathbbm{D}\backslash$ A at $\mathbbm{S}^{1}\setminus A$. Combined with the conformal invariance of the Brownian motion, it then implies that $g(\widetilde{\ell}_{1})$ is independent of the event $B[0,\tau_{\mathbbm{D}}]\cap A=\emptyset$. By varying $A$, we conclude that $g(\widetilde{\ell}_{1})$ is independent of the outer boundary of $B[0,\tau_{\mathbbm{D}}]$, and hence $\ell_{1}$. ∎
@@
In this section, we prove Theorem 1.1 based on the exact laws of the conformal radii in Section 3. Recall that $\mathsf{Bac}_{\varepsilon}$ is the event that there exist two disjoint subpaths on $B[0,\tau_{\mathbbm{D}}]$ from $\varepsilon\mathbbm{S}^{1}$ to $\frac{1}{2}\mathbbm{S}^{1}$. The lower bound of $\mathbbm{P}[\mathsf{Bac}_{\varepsilon}]$ in (1.2) then relies on the following continuum version of Menger’s theorem on the connectedness of a graph.
@@
$\mathsf{Bac}_{\varepsilon}$ occurs if $\ell_{1}\cap\mathbbm{A}_{1/2}\neq\emptyset$ and $\widetilde{\ell}_{1}\cap\varepsilon\mathbbm{D}\neq\emptyset$.
@@
Note that $B[s_{1},t_{1}]$ forms a connected, locally connected, and locally compact metric space when equipped with the Euclidean metric on $\mathbbm{C}$. By Definition 3.1, $B[s_{1},t_{1}]$ cannot be disconnected by removing any single point (otherwise it would contradict the definition of $s_{1}$). Furthermore, on the event $\{\ell_{1}\cap\mathbbm{A}_{1/2}\neq\emptyset\}\cap\{\widetilde{\ell}_{1}\cap\varepsilon\mathbbm{D}\neq\emptyset\}$, there are $s,t\in[s_{1},t_{1}]$ such that $B_{s}\in\varepsilon\mathbbm{D}$ and $B_{t}\in\mathbbm{A}_{1/2}$. Thus by Theorem 4.1, there exist two arcs on $B[s_{1},t_{1}]$ from $B_{s}$ to $B_{t}$, staying disjoint apart from their endpoints. The result then follows by taking the sub-trajectories of these two arcs from their last hitting points of $\varepsilon\mathbbm{S}^{1}$ to their first hitting points of $\frac{1}{2}\mathbbm{S}^{1}$, respectively. ∎
@@
By Lemma 4.2, we have $\mathbbm{P}[\mathsf{Bac}_{\varepsilon}]\geq\mathbbm{P}[\ell_{1}\cap\mathbbm{A}_{1/2}\neq\emptyset,\ \widetilde{\ell}_{1}\cap\varepsilon\mathbbm{D}\neq\emptyset]$. Note that $\mathbbm{P}[\ell_{1}\cap\mathbbm{A}_{1/2}\neq\emptyset]=p_{0}>0$, which does not depend on $\varepsilon$. On the other hand,
$\mathbbm{P}[\widetilde{\ell}_{1}\cap\varepsilon\mathbbm{D}\neq\emptyset|\ell_{1}\cap\mathbbm{A}_{1/2}\neq\emptyset]$ $\geq\mathbbm{P}[\mathrm{CR}(\widetilde{\ell}_{1},0)<\varepsilon|\ell_{1}\cap\mathbbm{A}_{1/2}\neq\emptyset]$
$\geq\mathbbm{P}[\mathrm{CR}(\rho,0)<\varepsilon|\ell_{1}\cap\mathbbm{A}_{1/2}\neq\emptyset]$
due to Schwartz’s lemma and that $\mathrm{CR}(\rho,0)=\frac{\mathrm{CR}(\widetilde{\ell}_{1},0)}{\mathrm{CR}(\ell_{1},0)}>\mathrm{CR}(\widetilde{\ell}_{1},0)$. By Lemma 3.5, $\mathrm{CR}(\rho,0)$ is independent of $\ell_{1}$. Therefore, combined with Corollary 3.7, we have
$\mathbbm{P}[\mathsf{Bac}_{\varepsilon}]\geq p_{0}\mathbbm{P}[\mathrm{CR}(\rho,0)<\varepsilon]\geq\frac{C_{1}}{\log|\log\varepsilon|}$
for some positive constant $C_{1}$, as desired. ∎
Now we turn to the upper bound of $\mathbbm{P}[\mathsf{Bac}_{\varepsilon}]$. We first need the following basic fact.
@@
Let $\varepsilon\in(0,10^{-2})$. Suppose $\ell\subset\mathbbm{D}$ is a simple loop such that $\mathrm{dist}(\ell,0)>5\sqrt{\varepsilon}$, and let $g:D(\ell)\to\mathbbm{D}$ be any conformal map such that $g(0)=0$. Then $\overline{g(\varepsilon\mathbbm{D})}\subset 5\sqrt{\varepsilon}\mathbbm{D}$ and $g(\frac{1}{2}\mathbbm{D}\cap D(\ell))\supset\frac{1}{2}\mathbbm{D}$.

### v2

@@ Abstract
Motivated by critical planar percolation, we investigate a “backbone” event of planar Brownian motion, i.e. the existence of two disjoint subpaths on the Brownian trajectory connecting the $\varepsilon$-neighborhood of the starting point to a macroscopic distance. We show that the probability of this event is $C(\log|\log\varepsilon|)^{-1}(1+o(1))$ as $\varepsilon\to 0$ for some constant $C\in(0,\infty)$.
@@ 1.1 Overview and the main result
Note that in the critical percolation, one can also define the monochromatic arm exponents corresponding to that there exist $n$ disjoint arms of the same color joining two boundaries of an annulus. Based on the relation between Brownian motion and critical percolation, it is natural to explore the analog of such monochromatic arm exponents in the context of Brownian motion. Recently, [NQSZ24] derives the exact value of percolation monochromatic 2-arm exponent, namely the backbone exponent. See references therein for more background on the backbone exponent of percolation.
In this paper, we investigate the Brownian counterpart, and find that such Brownian backbone probability indeed has an iterated logarithmic decay (so the “Brownian backbone exponent” is equal to 0). This also shows a big difference between Brownian motion and critical percolation, although their outer boundaries (or hulls) are the same.
@@
There exists a constant $C\in(0,\infty)$ such that,
$\lim\limits_{\varepsilon\to 0}\mathbbm{P}[\mathsf{Bac}_{\varepsilon}]\cdot(\log|\log\varepsilon|)=C.$ (1.2)
In fact, our proof gives an explicit expression of the constant $C$ in (1.2) via the quantity $S$ defined in (4.1); see (4.5) in Section 4.
@@
The same result also holds when $\frac{1}{2}\mathbbm{S}^{1}$ in (1.1) is changed to any fixed $r\mathbbm{S}^{1}$ with $r\in(0,1)$ (and for sufficiently small $\varepsilon$), except that the corresponding constant in (1.2) will depend on $r$.
Note that the backbone event $\mathsf{Bac}_{\varepsilon}$ is closely related to the spatial distribution of the cut points of $B[0,\tau_{\mathbbm{D}}]$. Indeed, our proof of Theorem 1.1 is based on our recent paper with Fu and Sun [CFSX25], from which we can define a layer structure for the cut points of $B[0,\tau_{\mathbbm{D}}]$ and then solve it explicitly; see Proposition 3.2 in Section 3. The remaining proof of Theorem 1.1 is then in a similar fashion to [JLQ23, Lemma 2.6]; see Section 4. We mention that [CFSX25] heavily relies on the connection between planar Brownian motion and $\operatorname{SLE}_{8/3}$, and especially their coupling with Liouville quantum gravity (LQG). It would be interesting to find a derivation of Theorem 1.1 without relying on LQG.
@@ 1.2 Outlook and discussions
  • One can also consider the backbone probability for three-dimensional Brownian motion, i.e. the probability that there are two disjoint subpaths joining $\varepsilon\mathbbm{S}^{2}$ and $\frac{1}{2}\mathbbm{S}^{2}$ on the Brownian tracjectory starting from 0 until hitting $\mathbbm{S}^{2}$. We conjecture that such probability equals $\varepsilon^{\alpha+o(1)}$ as $\varepsilon\to 0$ for some $\alpha\in(0,\infty)$. (The backbone probability for $d\geq 4$ dimensional Brownian motion becomes less interesting since the Brownian motion is now a simple curve.)
  • In our forthcoming work [CX26], we will extend the result of this paper to the Brownian loop soup cases. Let $(B_{t})_{0\leq t\leq\tau_{\mathbbm{D}}}$ be as before, and let $\mathcal{L}$ be an independent Brownian loop soup on $\mathbbm{D}$ with intensity $\frac{c}{2}$ for $c\in(0,1]$. For a subset $A\subset\overline{\mathbbm{D}}$, denote $\mathcal{C}(A)$ to be the union of $A$ and all loop-soup clusters in $\mathcal{L}$ intersecting with $A$, and let $H:=\mathcal{C}(B[0,\tau_{\mathbbm{D}}])$. However, unlike the Brownian motion case, there is a positive probability that 0 itself can have two disjoint subpaths (except on 0) on $H$ connected to $\frac{1}{2}\mathbbm{S}^{1}$. Moreover, in [CX26] we will show the following
@@
Organization of the paper. In Section 2, we first review the explicit relation between the laws of conformal radii and moduli for $\operatorname{SLE}_{8/3}$-type loops, which is an important ingredient of Theorem 1.1. Such relation was implicitly established in [ARS25b], and we will provide its proof in Appendix A for completeness. Then Sections 3 and 4 are devoted to prove Theorem 1.1.
@@
Acknowledgment. We are grateful to Xin Sun for many helpful comments and suggestions on the early draft of this paper. We also thank Xinyi Li for helpful discussions. G.C. and Z.X. were partially supported by National Key R&D Program of China (No. 2023YFA1010700). G.C. was partially supported by National Key R&D Program of China (No. 2021YFA1002700).
@@ 2 Conformal radii of $\operatorname{SLE}_{8/3}$-type loops
3 Exact laws of the conformal radii in $B[0,\tau_{\mathbbm{D}}]$
@@
Let $(t_{i})_{i\geq 1}$ be the collection of cut time $t$’s for $B[0,\tau_{\mathbbm{D}}]$ such that the outer boundary of $B[0,t]$ is a simple loop, ranked in decreasing order. Note that a.s. $\lim_{i\to\infty}t_{i}=0$. For each $i\geq 1$, define $s_{i}$ to be the last cut time before $t_{i}$ (thus $t_{i}>s_{i}>t_{i+1}$). Let $\ell_{i}$ and $\widetilde{\ell}_{i}$ be the outer boundary of $B[0,t_{i}]$ and the boundary of the connected component of $\mathbbm{D}\setminus B[s_{i},t_{i}]$ containing the origin, respectively.
@@
    Let $g_{n}:D(\ell_{n})\to\mathbbm{D}$ be the conformal map such that $g_{n}(0)=0$ and $g_{n}(B_{t_{n}})=1$. Then $g_{n}(\widetilde{\ell}_{n})$ is independent of $\ell_{n}$ and $\widetilde{\ell}_{n-1}$. In particular, $\frac{\mathrm{CR}(\widetilde{\ell}_{n},0)}{\mathrm{CR}(\ell_{n},0)}$ is independent of $\ell_{n}$ and $\widetilde{\ell}_{n-1}$.
@@
    Let $\widetilde{g}_{n}:D(\widetilde{\ell}_{n})\to\mathbbm{D}$ be the conformal map such that $\widetilde{g}_{n}(0)=0$ and $\widetilde{g}_{n}(B_{s_{n}})=1$. Then $(\widetilde{g}_{n}(B_{t}))_{t\in[0,s_{n}]}$ is independent of $(B_{t})_{t\in[s_{n},\tau_{\mathbbm{D}}]}$, and has the same law (up to a time change) as $(B_{t})_{t\in[0,\tau_{\mathbbm{D}}]}$ conditioned on $B_{\tau_{\mathbbm{D}}}=1$.
As a corollary, $\left(\frac{\mathrm{CR}(\widetilde{\ell}_{n},0)}{\mathrm{CR}(\ell_{n},0)}\right)_{n\geq 1}$ and $\left(\frac{\mathrm{CR}(\ell_{n+1},0)}{\mathrm{CR}(\widetilde{\ell}_{n},0)}\right)_{n\geq 0}$ are two i.i.d. r.v. sequences, and mutually independent of each other.
@@
We first prove (ii). Due to rotational invariance, it suffices to consider the case when conditioning on $B_{\tau_{\mathbbm{D}}}=1$. Let $(Y_{t})_{t\geq 0}$ be the time reversal of $(B_{t})_{t\in[0,\tau_{\mathbbm{D}}]}$ conditioned on $B_{\tau_{\mathbbm{D}}}=1$. Note that $Y_{t}$ is indeed a Brownian excursion in $\mathbbm{D}$ from 1 to 0. Let $(\mathcal{F}_{t})_{t\geq 0}$ be the natural filtration of $(Y_{t})_{t\geq 0}$, and $(\mathcal{G}_{t})_{t\geq 0}$ be the cut time filtration defined in [Vir03, Section 3] (i.e. $\mathcal{G}_{t}$ is the $\sigma$-field generated by $\mathcal{F}_{t}$ and the set of all cut times of $(Y_{t})_{t\geq 0}$ before time $t$). By [Vir03, Proposition 14], if $\tau$ is a $(\mathcal{G}_{t})_{t\geq 0}$-stopping time such that $\tau$ is also a cut time of $(Y_{t})_{t\geq 0}$, then the conditional law of $(Y_{\tau+u})_{u\geq 0}$ given $\mathcal{G}_{\tau}$ is a Brownian excursion from $Y_{\tau}$ to 0 in the connected component of $\mathbbm{D}\backslash Y[0,\tau]$ containing 0. Now we consider $s_{n}^{\prime}:=\tau_{\mathbbm{D}}-s_{n}$. Note that $s_{n}^{\prime}$ is the first cut time of $(Y_{t})_{t\geq 0}$ after $s_{n-1}^{\prime}$ such that $Y[s^{\prime}_{n-1},s^{\prime}_{n}]$ disconnects the origin from $\mathbbm{S}^{1}$, thus is a $(\mathcal{G}_{t})_{t\geq 0}$-stopping time. Therefore, the conditional law of $(\widetilde{g}_{n}(Y_{s_{n}^{\prime}+u}))_{u\geq 0}$ given $\mathcal{G}_{s_{n}^{\prime}}$ is an independent Brownian excursion (up to a time change) in $\mathbbm{D}$ from $1$ to 0. By taking the time reversal of $(Y_{t})_{t\geq 0}$, we obtain the result.
Then we prove (i). According to (ii), $(\widetilde{g}_{n-1}(B_{t}))_{t\in[0,s_{n-1}]}$ is independent of $(\widetilde{\ell}_{n-1},B_{s_{n-1}})$, and has the same law (up to a time change) as $(B_{t})_{t\in[0,\tau_{\mathbbm{D}}]}$ conditioned on $B_{\tau_{\mathbbm{D}}}=1$. Since $(g_{n}(\widetilde{\ell}_{n}),\widetilde{g}_{n-1}(\ell_{n}))$ is determined by $(\widetilde{g}_{n-1}(B_{t}))_{t\in[0,s_{n-1}]}$, it follows that $(g_{n}(\widetilde{\ell}_{n}),\widetilde{g}_{n-1}(\ell_{n}))$ is independent of $(\widetilde{\ell}_{n-1},B_{s_{n-1}})$.
[…]
Combining the previous two paragraphs, we see that $g_{n}(\widetilde{\ell}_{n})$, $\widetilde{g}_{n-1}(\ell_{n})$, $(\widetilde{\ell}_{n-1},B_{s_{n-1}})$ are mutually independent. Since $\widetilde{g}_{n-1}$ is determined by $\widetilde{\ell}_{n-1}$ and $B_{s_{n-1}}$, we have $\sigma(\widetilde{g}_{n-1}(\ell_{n}),\widetilde{\ell}_{n-1},B_{s_{n-1}})$ equals $\sigma(\ell_{n},\widetilde{\ell}_{n-1},B_{s_{n-1}})$. Therefore, $g_{n}(\widetilde{\ell}_{n})$ is independent of the triple $(\ell_{n},\widetilde{\ell}_{n-1},B_{s_{n-1}})$, as desired. ∎
@@
In this section, we prove Theorem 1.1 based on the exact laws of the conformal radii in Section 3. The general proof steps are akin to [JLQ23, proof of Lemma 2.6]. Recall that $\mathsf{Bac}_{\varepsilon}$ is the event that there exist two disjoint subpaths on $B[0,\tau_{\mathbbm{D}}]$ from $\varepsilon\mathbbm{S}^{1}$ to $\frac{1}{2}\mathbbm{S}^{1}$. We first need the following continuous version of Menger’s theorem on the connectedness of a graph.
@@
$\mathsf{Bac}_{\varepsilon}$ occurs if $\ell_{n}\cap\mathbbm{A}_{1/2}\neq\emptyset$ and $\widetilde{\ell}_{n}\cap\varepsilon\mathbbm{D}\neq\emptyset$ for some $n\geq 1$.
@@
Recall $s_{n}$ and $t_{n}$ in Definition 3.1. Note that $B[s_{n},t_{n}]$ forms a connected, locally connected, and locally compact metric space when equipped with the Euclidean metric on $\mathbbm{C}$, and $B[s_{n},t_{n}]$ cannot be disconnected by removing any single point (otherwise it would contradict the definition of $s_{n}$). Furthermore, on the event $\{\ell_{n}\cap\mathbbm{A}_{1/2}\neq\emptyset\}\cap\{\widetilde{\ell}_{n}\cap\varepsilon\mathbbm{D}\neq\emptyset\}$, there are $s,t\in[s_{n},t_{n}]$ such that $B_{s}\in\varepsilon\mathbbm{D}$ and $B_{t}\in\mathbbm{A}_{1/2}$. Thus by Theorem 4.1, there exist two arcs on $B[s_{n},t_{n}]$ from $B_{s}$ to $B_{t}$, staying disjoint apart from their endpoints. The result then follows by taking the sub-trajectories of these two arcs from their last hitting points of $\varepsilon\mathbbm{S}^{1}$ to their first hitting points of $\frac{1}{2}\mathbbm{S}^{1}$, respectively. ∎
We also need the following input.
Lemma 4.3.
Define
$S:=\sum_{n\geq 1}\mathbbm{P}[\ell_{n}\cap\mathbbm{A}_{1/2}\neq\emptyset].$ (4.1)
Then $S\in(0,\infty)$.
Proof.
For $n\geq 2$, we have
$\mathbbm{P}[\ell_{n}\cap\mathbbm{A}_{1/2}\neq\emptyset]$ $\leq\mathbbm{P}[\mathrm{CR}(\ell_{n},0)\geq e^{-(n-1)}]+\mathbbm{P}[\ell_{n}\cap\mathbbm{A}_{1/2}\neq\emptyset,\mathrm{CR}(\ell_{n},0)<e^{-(n-1)}]$
$\leq\mathbbm{P}[n-1\geq\log\mathrm{CR}(\ell_{n},0)^{-1}]+\mathbbm{P}[\ell_{n}\cap\mathbbm{A}_{1/2}\neq\emptyset,\mathrm{dist}(\ell_{n},0)<e^{-(n-1)}]$
$\leq\mathbbm{P}\left[n-1\geq\sum_{i=1}^{n-1}\log\frac{\mathrm{CR}(\widetilde{\ell}_{i-1},0)}{\mathrm{CR}(\widetilde{\ell}_{i},0)}\right]+\mathbbm{P}[\ell_{n}\cap\mathbbm{A}_{1/2}\neq\emptyset,\mathrm{dist}(\ell_{n},0)<e^{-(n-1)}].$ (4.2)
By Lemma 3.5, $\left(\log\frac{\mathrm{CR}(\widetilde{\ell}_{i-1},0)}{\mathrm{CR}(\widetilde{\ell}_{i},0)}\right)_{i\geq 1}$ are i.i.d. random variables. Furthermore, Corollary 3.7 yields that $\mathbbm{P}\left[\log\frac{1}{\mathrm{CR}(\widetilde{\ell}_{1},0)}>t\right]=\frac{3}{\log t}(1+o(1))$ as $t\to\infty$, hence $\mathbbm{E}\left[\log\frac{1}{\mathrm{CR}(\widetilde{\ell}_{1},0)}\right]=+\infty$. By the large deviation principle, there exists $u\in(0,1)$ such that
$\mathbbm{P}\left[n-1\geq\sum_{i=1}^{n-1}\log\frac{\mathrm{CR}(\widetilde{\ell}_{i-1},0)}{\mathrm{CR}(\widetilde{\ell}_{i},0)}\right]\leq u^{n-1}.$ (4.3)
We now bound the second term on the right side of (4.2) by the Brownian disconnection exponent derived in [LSW01b]. For $r\in(0,1)$, let $\tau_{r}=\inf\{t\geq 0:|B_{t}|=r\}$ and $\sigma_{r}=\sup\{t\in[0,\tau_{\mathbbm{D}}]:|B_{t}|=r\}$. Then for $s<r<1$, $(B_{t})_{\sigma_{s}\leq t\leq\tau_{r}}$ has the same law as the Brownian excursion on annulus $\mathbbm{A}_{s}\backslash\mathbbm{A}_{r}$. By [LSW01b], the probability that $(B_{t})_{\sigma_{s}\leq t\leq\tau_{r}}$ does not disconnect $s\mathbbm{S}^{1}$ and $r\mathbbm{S}^{1}$ equals $(\frac{s}{r})^{1/4+o(1)}$ as $s\to 0$ (we remark that such probability was exactly computed in the recent paper [CFSX25]). Since $(\ell_{n}\cap\mathbbm{A}_{1/2}\neq\emptyset,\mathrm{dist}(\ell_{n},0)<e^{-(n-1)})$ implies that $B[\sigma_{e^{-(n-1)}},\tau_{1/2}]$ does not disconnect $e^{-(n-1)}\mathbbm{S}^{1}$ and $\frac{1}{2}\mathbbm{S}^{1}$, we find
$\mathbbm{P}[\ell_{n}\cap\mathbbm{A}_{1/2}\neq\emptyset,\mathrm{dist}(\ell_{n},0)<e^{-(n-1)}]\leq e^{-\frac{1}{4}(n-1)(1+o(1))}$ (4.4)
as $n\to\infty$. The result then follows by combining (4.2), (4.3) and (4.4). ∎
In the following, we will prove Theorem 1.1 by showing that
$\lim_{\varepsilon\to 0}\mathbbm{P}[\mathsf{Bac}_{\varepsilon}]\cdot(\log|\log\varepsilon|)=3S,$ (4.5)
where $S$ is defined in (4.1). We first give the lower bound.
@@
For $n\geq 1$, let $A_{n,\varepsilon}=\{\ell_{n}\cap\mathbbm{A}_{1/2}\neq\emptyset,\widetilde{\ell}_{n}\cap\varepsilon\mathbbm{D}\neq\emptyset\}$, which satisfies $\cup_{n\geq 1}A_{n,\varepsilon}\subset\mathsf{Bac}_{\varepsilon}$ by Lemma 4.2. Furthermore, let $B_{n,\varepsilon}\subset A_{n,\varepsilon}$ be the event $\{\mathrm{dist}(\widetilde{\ell}_{n-1},0)>\varepsilon,\ell_{n}\cap\mathbbm{A}_{1/2}\neq\emptyset,\frac{\mathrm{CR}(\widetilde{\ell}_{n},0)}{\mathrm{CR}({\ell_{n}},0)}\leq\varepsilon\}$. Note that $B_{n,\varepsilon}\cap B_{m,\varepsilon}=\emptyset$ for every $n\neq m$. Thus, we have
$\mathbbm{P}[\mathsf{Bac}_{\varepsilon}]\geq\sum_{n\geq 1}\mathbbm{P}[B_{n,\varepsilon}]=\mathbbm{P}[\mathrm{CR}(\rho,0)<\varepsilon]\cdot\sum_{n\geq 1}\mathbbm{P}[\mathrm{dist}(\widetilde{\ell}_{n-1},0)>\varepsilon,\ell_{n}\cap\mathbbm{A}_{1/2}\neq\emptyset],$ (4.6)
where we use Lemma 3.5 that $\frac{\mathrm{CR}(\widetilde{\ell}_{n},0)}{\mathrm{CR}({\ell_{n}},0)}$ is independent of $\ell_{n}$ and $\widetilde{\ell}_{n-1}$ in the second equality (recall that $\rho$ is the loop defined above Proposition 3.2). Combining (4.6) with Corollary 3.7 and Lemma 4.3, we obtain that $\liminf_{\varepsilon\to 0}\mathbbm{P}[\mathsf{Bac}_{\varepsilon}]\cdot(\log|\log\varepsilon|)\geq 3S$, as desired. ∎
Now we turn to the upper bound of Theorem 1.1, i.e. $\limsup_{\varepsilon\to 0}\mathbbm{P}[\mathsf{Bac}_{\varepsilon}]\cdot(\log|\log\varepsilon|)\leq 3S$. This is based on the following basic fact.
@@
Let $\varepsilon\in(0,10^{-2})$. Suppose $\ell\subset\mathbbm{D}$ is a simple loop surrounding the origin with $\mathrm{dist}(\ell,0)>2\sqrt{\varepsilon}$, and let $f:D(\ell)\to\mathbbm{D}$ be any conformal map such that $f(0)=0$. Then $f(\varepsilon\mathbbm{D})\subset 2\sqrt{\varepsilon}\mathbbm{D}$.